OpenStudy (anonymous):

simplify: 4 to the third power(square root)3x+5 to the third power(square root)10x

OpenStudy (anonymous):

Raise both sides to 2 to cancel the square root sign. 4+3x=5^2, 4+3x=25, 3x=25-4, 3x=21. Therefore x=7 I think......

OpenStudy (anonymous):

sorry i typed that in the wrong place